### v1

# Enhancing Arabic Automated Essay Scoring with Synthetic Data and Error Injection

Abstract
[…]
utf8
Enhancing Arabic Automated Essay Scoring
with Synthetic Data and Error Injection
Chatrine Qwaider,1 Bashar Alhafni,1,2 Kirill Chirkunov,1 Nizar Habash,1,2 Ted Briscoe1 1MBZUAI, 2New York University Abu Dhabi {chatrine.qwaider,kirill.chirkunov,ted.briscoe}@mbzuai.ac.ae {alhafni,nizar.habash}@nyu.edu
[…]
utf8
Enhancing Arabic Automated Essay Scoring
with Synthetic Data and Error Injection
Chatrine Qwaider,1 Bashar Alhafni,1,2 Kirill Chirkunov,1 Nizar Habash,1,2 Ted Briscoe1 1MBZUAI, 2New York University Abu Dhabi {chatrine.qwaider,kirill.chirkunov,ted.briscoe}@mbzuai.ac.ae {alhafni,nizar.habash}@nyu.edu
1 Introduction
Automated Essay Scoring (AES) is a technology that automates the evaluation and scoring of essays to assess language learners’ writing quality while eliminating the need for human intervention Shermis and Burstein (2003). AES has gained great interest due to its significant benefits in the field of education Lagakis and Demetriadis (2021); Susanti et al. (2023). AES systems help teachers evaluate many essays with consistent scoring and reduced workload. On the other hand, AES helps students improve their writing quality through rapid real-time scoring and feedback Hahn et al. (2021). Unlike for English, for Modern Standard Arabic (MSA) it is difficult to develop robust and scalable AES systems, primarily due to the lack of essay datasets necessary for building effective Arabic AES Lim et al. (2021); Elhaddadi et al. (2024). This paper presents our proposed framework to tackle the issue of data scarcity and quality by utilizing Transformers and Large Language Models (LLMs) to generate and build a synthetic dataset. Our approach begins with prompting GPT-4o to generate a variety of Arabic essays covering multiple topics and different writing proficiency levels as defined by the Common European Framework of Reference (CEFR) Council of Europe (2001). Subsequently, we use a controlled error injection model to introduce errors into the correct Arabic essays, ensuring that erroneous essays reflect the type of errors that are commonly made by learners of Arabic in real-world scenarios. Our error injection approach consists of two steps: (i) Error Type Prediction, where a fine-tuned CAMeLBERT MSA model Inoue et al. (2021) classifies the most likely error type for each word, and (ii) Error Realization, where we apply a bigram MLE model to determine the most probable transformation for each predicted error type. Our framework enables the generation of realistic human-like essays, enhancing data augmentation for Arabic AES systems. Our main contributions are as follows:
  • Proposing a framework based on LLMs and Transformers for augmenting Arabic essays that accurately reflect human writing patterns.
  • Creating a synthetic Arabic AES dataset with 3,040 essays annotated with CEFR proficiency levels. [^1] The full dataset will be publicly available for research.
  • Developing an Arabic essay auto-marking system using a BERT-based scoring model, enabling accurate and scalable evaluation of Arabic essays based on CEFR standards.
The rest of the paper is organized as follows: §2 reviews related work on AES, §3 describes the dataset, and §4 outlines our data augmentation approach. §5 details the error injection methods, followed by evaluation in §6 and §7 presents conclusion and future work.
2 Related Work
The field of AES has been investigated extensively, particularly in English Lim et al. (2021); Ramesh and Sanampudi (2022), where some valuable tools have been introduced such as IntelliMetric Elliott et al. (2003), e-rater Attali and Burstein (2006), Grammarly, [^2] https://app.grammarly.com/ Write and Improve [^3] https://writeandimprove.com/ Yannakoudakis et al. (2018), and others. The development and validation of English AES have been enabled by the large-scale annotated datasets that have been released, such as the First Cambridge English (FCE) dataset [^4] https://ilexir.co.uk/datasets/index.html Yannakoudakis et al. (2011), Automated Student Assessment Prize (ASAP) dataset, [^5] https://www.kaggle.com/c/asap-aes the TOEFL11 corpus Blanchard (2013), and the ICLE (International Corpus of Learner English) Granger (2003). These datasets contained thousands of student essays with proficiency-level grades, often along multiple dimensions. In contrast, research on Arabic AES has not received comparable attention. Some studies have focused on Arabic AES, utilizing feature engineering and applying machine learning to develop models Alghamdi et al. (2014); Al-Shalabi (2016); Alobed et al. (2021); Gaheen et al. (2021). This work and related research has only partially addressed the key challenges hindering the development of Arabic AES models, namely, the lack of publicly available large-scale annotated datasets designed to improve the quality of Arabic writing. Ghazawi and Simpson (2024) introduced AR-AES, a benchmark dataset that contains 2,046 essays written by undergraduate students across three different university faculties, featuring a mix of question types. Two educators from each faculty annotated and graded the essays using clearly defined rubrics, aiming to assess students’ academic performance across various disciplines. Unlike AR-AES, our work aims to evaluate and assess writing proficiency using the CEFR rubrics as these constitute an internationally accepted standard. QAES Bashendy et al. (2024) presented the first publicly available trait-specific annotations for Arabic AES. QAES is based on the work of the Qatari Corpus of Argumentative Writing (QCAW) Ahmed et al. (2024), which consists of 195 Arabic argumentative essays. They proposed a multi-layer of annotation traints such as coherence, organization, grammar, and others. Despite its excellent annotation, it is small in size and limited to two prompts. While QAES multi-traits scores are publicly available, the QCAW holistic score is not public. Habash and Palfreyman (2022) presented the Zayed University Arabic-English Bilingual Undergraduate Corpus (ZAEBUC). This corpus comprises non-parallel essays in Arabic and English related to three prompts collected from first-year university students with differing writing proficiency. ZAEBUC includes 216 annotated Arabic essays featuring manual annotations for syntactic and morphological characteristics and a CEFR-based proficiency assessment. Again, ZAEBUC is small in size and limited to three prompts. Researchers have investigated several approaches to data augmentation such as sampling techniques, introducing noise into texts, or paraphrasing, to overcome the challenges of data scarcity and quality Li et al. (2022). The recent development of LLMs has paved the way for researchers to explore promising new data synthesis solutions Wang et al. (2024a); Long et al. (2024). Transformers and LLMs can closely mirror real-world distributions while introducing valuable variations across multiple tasks and domains Wang et al. (2024b). GPT models such as ChatGPT, GPT-4, and GPT-4o demonstrate significant capabilities in augmenting and generating synthetic essays for the purpose of English AES Ramesh and Sanampudi (2022). LLMs and Transformer models with Arabic capabilities have shown robust generalization across different Arabic NLP tasks such as: Question-Answering Samuel et al. (2024), Code-Switching Alharbi et al. (2024), Name Entity Recognition (NER) Sabty et al. (2021), Grammar Error Correction (GEC) Solyman et al. (2023), and Sentiment Analysis Refai et al. (2023). However, to the best of our knowledge no research has utilized such models to generate Arabic essays across CEFR writing proficiency levels.
\<وسائل التواصل الاجتماعي وتأثيرها على الفرد والمجتمع.>
How do social media affect individuals and society?
\<كيف نعزز ثقافة التسامح في المجتمع؟>
How can the UAE promote a culture of tolerance in society?
\<التطور الحضاري الذي تشهده دولة الإمارات العربية المتحدة>
What do you think are the most important developments in the UAE at the moment?
Table 1: The prompts given to the essay writers in the ZAEBUC corpus Habash and Palfreyman (2022).
3 Data: The ZAEBUC Corpus
For all our experiments and reported results, we use the ZAEBUC corpus Habash and Palfreyman (2022). ZAEBUC comprises essays written by native Arabic speakers, which were manually corrected and annotated for writing proficiency using the CEFR Council of Europe (2001) rubrics and scale. Each essay was independently annotated by three CEFR-proficient bilingual speakers. Habash and Palfreyman (2022), assigned a holistic CEFR level to each essay by converting the three CEFR ratings into numerical scores (ranging from 1 to 6) and then taking the rounded average. The essays in the corpus were limited to three prompt choices on Social Media, Tolerance, and Development; see Table 1. We use the splits created by Alhafni et al. (2023). Table 2 presents the ZAEBUC corpus CEFR level distributions in terms of the holistic CEFR levels.
CEFR Level Count Percentage
A1 0 0%
A2 7 3%
B1 110 51%
B2 80 37%
C1 11 5%
C2 0 0%
Unassessable 6 3%
Total 214 100%
Table 2: ZAEBUC corpus CEFR level distributions.
4 Synthetic Data Augmentation
We propose a synthetic data augmentation approach leveraging the ZAEBUC dataset to generate synthetic essays that align with CEFR rubrics and have features similar to human text. The pipeline utilizes three phases: Building Essay Prompts, Feature Profiling, and finally Data Augmentation.
4.1 Building Essay Prompts
We began by compiling a collection of essay prompts covering diverse categories and proficiency levels. We considered three proficiency levels aligned with the CEFR scale: Beginner (A1–A2), Intermediate (B1–B2), and Advanced (C1–C2). While some categories, such as Hobbies, are suitable for all levels, others, like Politics, Technology, and Education, were introduced specifically for advanced learners. Using various LLMs, including GPT-4o, Gemini, and Copilot, we generated 100 prompts across these categories, ensuring relevance for Arabic essay writing and an unbiased distribution across proficiency levels. We then manually reviewed the prompts, removing redundancies and irrelevant entries while maintaining comprehensive coverage. The final collection consists of 152 balanced and diverse prompts. Table 3 presents the selected categories and the distribution of prompts across levels, while Table 4 provides example prompts from the Hobbies category.
Topic B I A
Culture and Traditions 1 3 2
Daily Life 2 2 2
Education 3 6 8
Environment 2 2 3
Future 1 2 2
History and Culture 2 2 2
Hobbies 3 2 2
Imaginary 5 2 2
Life/Time Management 4 4 2
Personal Experiences 7 2 2
Relations 4 2 2
School Life 4 2 2
Sport and Health 2 3 1
Technology and Media 2 8 6
Travel and Experience 1 2 1
Politics and Government 2 2 7
Social Issues 2 7 6
Total 47 53 52
Table 3: Count of Arabic text prompts by level and topic. B: Beginner level (A1, A2), I: Intermediate level (B1, B2), A: Advanced level (C1, C2).
Level Arabic Prompt English Prompt
Beginner \<ما هي هوايتك المفضلة؟ >  $\bullet$ $\bullet$ What is your favorite hobby?
\<تحدث عن نشاط تحبه في عطلة نهاية الأسبوع.>  $\bullet$ $\bullet$ Talk about a weekend activity you love.
\< ما هي الرياضة التي تحب ممارستها؟ لماذا؟>  $\bullet$ $\bullet$ What is your favorite sport? Why?
Intermediate \<في ممارستها؟ لماذا تحبها؟> \<ما هي هوايتك المفضلة وكيف بدأت >   $\bullet$ $\bullet$ What is your favorite hobby and how did you start practicing it? Why do you enjoy it?
\< ما هي ولماذا تهمك؟> \< هل ترغب في تعلم هواية جديدة؟>   $\bullet$ $\bullet$ Do you wish to learn a new hobby? What is it and why does it interest you?
Advanced \<ناقش كيف تؤثر الهواية على صحتك النفسية والجسدية.>   $\bullet$ $\bullet$ Discuss how hobbies impact your mental and physical health.
\<كيف يمكن أن تكون الهوايات وسيلة للتعبير عن الذات؟> How can hobbies serve as a means of self-expression?
\< هل هناك هواية جديدة ترغب في تجربتها؟ >   $\bullet$ $\bullet$ Is there a new hobby you wish to try? Discuss the reasons
\<ناقش الأسباب التي تجعلك مهتمًا بها وكيف تعتقد أنها ستفيدك.> you are interested in it and how you believe it will benefit you.
Table 4: Examples of prompts related to the topic of Hobbies and classified into one of three different levels.
4.2 Feature Profiling
We construct linguistic profiles for each CEFR level using the ZAEBUC corpus. Each profile contains various levels of linguistic information. Representing different lexical and syntactic features, we use the number of words/sentences ($N_{w}$,$N_{s}$), the number of tokens /vocabulary ($N_{v}$), words/sentences lengths ($L_{w}$,$L_{s}$), and sentence complexity measured by syntactic tree depth ($D_{s}$). We define the lexical diversity (Type-Token Ratio, TTR) as:
TTR=Unique TokensTotal $\text{TTR}=\frac{\text{Unique Tokens}}{\text{Total Tokens}}$ (1)
Similarly, we calculate the sentence complexity by:
$C_{s}=\frac{\sum_{i=1}^{N}D_{i}}{N}$ (2)
where $D_{i}$ is the syntactic depth of sentence $i$ and $N$ is the total number of sentences. For morphological features, we use the ZAEBUC morphological annotations: the most frequent POS tags, such as nouns, verbs, adjectives, etc. We aggregate all extracted features across the essays to get a quantitative representation at different writing CEFR levels, which serves as a reference for later stages.
4.3 Zero-shot Data Augmentation
Effective LLM prompt engineering has become increasingly important, as the model’s output varies based on the prompt, provided instructions, and prompt language. Previous studies in Arabic NLP have shown that using English as the instruction language for input prompts can improve output quality Kmainasi et al. (2024); Koto et al. (2024). In our approach, we experiment with various prompts for zero-shot data augmentation to identify those that produce human-like text while adhering to guideline instructions. We use GPT-4o as our generation model due to its affordability and larger token capacity for both input and output. The GPT prompts include: (a) the target CEFR level, (b) CEFR guidelines and instructions, (c) the linguistic profile for the targeted CEFR level to control the prompt output, and (d) the topic prompt or question from the previously mentioned topic prompts dataset. To check the quality of the generated essays and whether they follow the prompt instructions, we build a linguistic feature profile for each augmented essay. We employ the cosine similarity to assess the faithfulness of the augmented approach and whether the augmented essays are close to the original CEFR profiles. We applied the cosine similarity as in equation 3 either to one of the features profiles (lexcial/syntactical features , morphological features or both of them).
$\cos(\theta)=\frac{\sum_{i}P_{i}Q_{i}}{\sqrt{\sum_{i}P_{i}^{2}}\cdot\sqrt{\sum_{i}Q_{i}^{2}}}$ (3)
where $P_{i}$ and $Q_{i}$ are the probability values of linguistic features in the original and generated distribution profiles. This metric ensures that the synthetic data closely aligns with real human essay patterns. Later we calculate the matching agreements by:
$\text{Agreement}=\frac{\sum_{i=1}^{n}(\hat{y}_{i}=y_{i})}{n}$ (4)
where $\hat{y}_{i}$ is the predicted level and $y_{i}$ is the ground truth. We conducted multiple rounds of prompt engineering refinements to improve the quality of the generated Arabic essays and ensure alignment with CEFR levels. First, we found that straightforward prompts without explicit controlled linguistic instructions and explanations resulted in incoherent essays, including irrelevant topics and English text, achieving only 20.5% matching agreement with linguistic feature profiles. In a subsequent round, we introduced detailed definitions of linguistic features and restricted outputs to Arabic-only text, which improved agreement to 26%. However, the model still occasionally produced incomplete essays and injected text from the prompt into the essay. The most effective prompt structure format is illustrated in Figure 1. We separated system-level control instructions from user-defined parameters, thereby providing clearer guidance for structured and proficiency-aligned text generation. This refinement increased agreement to 27.5%, demonstrating that precisely controlled instructions enhance LLM performance in structured writing tasks. Ultimately, we generated 3,040 Arabic essays covering all CEFR levels and various topics, where each prompt produced ten essays. The structured and controlled prompt design not only improved alignment with learner writing styles but also established a consistent framework for synthetic Arabic text generation. Table 5 presents the distribution of generated essays across different CEFR levels.
CEFR Level Count Percentage
A1 470 15.5%
A2 470 15.5%
B1 530 17.4%
B2 530 17.4%
C1 520 17.1%
C2 520 17.1%
Total 3,040 100%
Table 5: The generated corpus CEFR level distributions.
Figure 1: GPT-4o prompts messages that has been used to generate Arabic essays
5 Error Injection
Human-generated text naturally contains some grammatical errors and linguistic infelicities. In order to generate human-like essays, we need to include similar kinds of errors that reflect the level of writing attainment. In this phase, we prompt GPT-4o to inject errors into the previously generated essays while maintaining their aligned CEFR levels by utilizing error profiling.
5.1 Error Profiling
Error Distribution Profiles
To model the distribution of errors to inject into the synthetic essays, we again leverage the ZAEBUC corpus, which contains both the erroneous essays aligned with the manually corrected ones. We followed the same methodology we used to construct the linguistic feature profiles for each CEFR level to develop error distribution profiles aligned with CEFR levels. The error profile captures and reflects the authentic distribution patterns observed in human writing at different CEFR levels. Developing an Error Instruction Repository To prompt GPT-4o to generate essays containing errors we applied the GED model proposed by Alhafni et al. (2023) to the ZAEBUC corpus to annotate errors using 13 error tags and to obtain error distributions for each CEFR level. We created the repository using the same 13 tags (ReplaceM, ReplaceO, ..), where we also added a formal definition of what those tags describe in terms of linguistic errors. In addition, we expanded the error taxonomy by splitting it into finer-grained classes. Each error instruction is followed by an example showing the correct word and the erroneous version. The explanation was based on the extended ALC taxonomy Alfaifi et al. (2013), which was refined later and introduced as ARETA Belkebir and Habash (2021). Figure 2 shows some examples from our error instruction repository.
Figure 2: an Example of orthographical error instructions from the developed errors instructions repository
5.2 GPT-Based Error Injection
We prompted GPT-4o to inject errors into the synthetically generated essays based on the error distribution profiles while maintaining the CEFR level. The model processed one essay at a time in a zero-shot setting, except that we included the definition and explanation of the error tags. For example, Replace_M indicates a morphological error, while Merge targets two mistakenly split tokens that need to be merged, and so on. After conducting multiple experiments, we observed the following issues: (i) The model struggled to follow the predefined error distribution perhaps due to the complexity of the prompts. (ii) The model was confused by certain error tags, particularly SPLIT, MERGE as well as the UNK (unknown errors). These errors were mainly ignored in the injected text. (iii) We calculated the cosine similarity between the main error profile and the injected essays’ error distributions as shown in equation 3. When we injected all errors at once, the similarity agreements did not exceed 20%; however, when we reduced the number of error tags per essay or omitted the misleading tags, the agreements significantly improved, reaching 86% Therefore we implemented a method where each error type was injected separately. This required multiple iterations over the same essay, corresponding to the number of error tags shown in the error distribution profile for each CEFR level. Figure 3 shows an example of a GPT-4o prompt for error injection. Some errors occurred more frequently than others among Arabic writers. For example, orthographic errors are more frequent than semantic errors. So, rather than repeating a single error instruction for all instances of an error tag, we randomly select a weighted number of error instructions proportional to the average occurrence of each error type. Figure 4 shows the pseudocode for the selection process. The full pseudocode is in Appendix A.
Figure 3: Sample GPT-4o error injection prompt
Figure 4: Error injection based on average error count
5.3 Controlled Error Injection
We introduce a controlled method for injecting errors into clean text, ensuring that the resulting erroneous sentences follow the empirical error distributions observed at each CEFR level. More formally, given an input sentence ($X$) and its CEFR level ($L$), we introduce errors in two steps: Error Type Prediction and Error Realization.
Error Type Prediction
We estimate the probability of an error type occurring at a given word, i.e., $P(error\_type|word)$. To do so, we leverage ARETA in a reverse annotation process where we process correct–erroneous sentence pairs, tagging each correct word with its corresponding error type. Using this annotated data, we train a token-level BERT classifier to predict the most likely error type for each word in a given correct sentence. We fine-tune CAMeLBERT MSA Inoue et al. (2021) to build our classifier.
Error Realization
To determine how a word should be corrupted, we first align correct–erroneous sentence pairs using the algorithm proposed by Alhafni et al. (2023). For each aligned pair, we extract edit transformations that capture the operations required to convert a correct word into its erroneous counterpart. Using this data, we estimate $P(transformation|error\_type)$ with a bigram Maximum Likelihood Estimation (MLE) lookup model: $count(transformation,error\_type)$ / $count(error\_type)$. During inference, we apply the BERT classifier to predict error types for each word in a sentence. We then filter these predictions, retaining only error types relevant to the sentence’s CEFR level. Finally, the MLE model selects the most probable corruption for a given error type. A complete example of a B1–level essay generated by the proposed model is in Figure 5.
Figure 5: An Example of a B1 Arabic Essay generated by GPT-4o using the Hobbies prompt and the same essay after injecting errors by the controlled BERT-based model.
Average Reference Multi-Reference
Train Data QWK Acc F1 P R QWK Acc F1 P R
Zaebuc-1 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
Zaebuc-1 + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
Zaebuc-1 + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
Zaebuc-1 + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
Zaebuc-1 + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
Table 6: Performance comparison of different training datasets. GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one shot error example, while GPT errors_0 with Zero shot settings
6 Evaluation
6.1 Metrics
Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely-used metric in recent AES research Ke and Ng (2019). We also report accuracy, macro precision (P), recall (R), and F1 scores. Model predictions are evaluated against the average of the gold holistic CEFR labels and in a multi-reference setting against the three individual gold labels per essay (§3).
6.2 Data
We use the ZAEBUC dataset for all the experiments. The dataset was split (70%,15%,15%) for Train, Dev and Test sets, respectively Alhafni et al. (2023). Table 6 shows the performance results of the BERT-based Scoring Model utilizing different training sets.
6.3 Model
We treat the task of automated essay scoring (AES) as a text classification problem. For Arabic AES, we fine-tune CAMeLBERT MSA Inoue et al. (2021) on the training split of ZAEBUC. The models were trained by using the average CEFR gold labels. During training, we ignore the essays that are labeled as Unassessable, but we penalize the models for missing them in the evaluation. We fine-tune the models for 5 epochs, with a maximum sequence length of 512, a learning rate of 5e-5, and a batch size of 32.
6.4 Results
Our results are presented in Table 6. Our baseline system, only trained on the ZAEBUC training set, indicates room for improvement, with the F1 at 24.50% and QWK at 22.44%. We then switched between different datasets to enhance the Arabic AES model and to measure the impact of data augmentation on the model. We introduced the 3,040 correct essays generated using GPT-4o, which dropped the QWK compared to the baseline while increasing the F1. However, we observed significant improvements in the multi-reference settings for the QWK (96.00%) and F1 (92.32%). Since this is an error-free, synthetic dataset, we further refine the model by introducing essays that mimic human-like writing styles. We compare the two approaches introduced in section 5: (1) GPT-based error injection (with and without error instruction examples) and (2) the controlled error injection method based on CAMeLBERT MSA. Inoue et al. (2021). The results demonstrate that the controlled error model enhances performance across all metrics, particularly in the average-reference setting, that achieved 27.87 % and 38.02 for QWK and F1 respectively. However, the GPT-based error injection performs best under the multi-reference setting, where the one-shot essays increased the QWK to 96.47 and the zero-shot erroneous essays improved the F1 to 95.12%. These results suggest that LLMs and Transformer methods generalize effectively in AES models and also improve performance. The results emphasize that a controlled methodology is more reliable in mimicking patterns of learner errors and mistakes. This positions the controlled BERT-based model as a vital component of the data augmentation pipeline, effectively addressing the limitations of the GPT models.
7 Conclusions and Future Work
This paper proposes a hybrid framework for building an Arabic AES system. We utilize both LLMs and Transformers to overcome the challenges of dataset scarcity by generating synthetic data that mimics the writing styles of Arabic learners. We leverage the ZAEBUC corpus and build linguistic feature profiles as well as error profiles that align with each CEFR score. We prompt GPT-4o to generate 3,040 Arabic essays based on multiple CEFR levels corresponding to 152 prompts. However, GPT-4o’s effectiveness depends heavily on prompt engineering, achieving only 27.5% agreement with the original linguistic feature profiles. To introduce errors into GPT-generated essays, we compare two approaches: (1) GPT-4o error injection, which struggles with realistic learner writing due to misinterpreted linguistic errors or complex prompts, causing inconsistencies; (2) a controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their natural occurrence. The latter method is an essential component of data augmentation for Arabic AES. Our evaluation, using a fine-tuned BERT classifier, shows that the proposed hybrid framework that combines GPT-generated data and a controlled error injection method shows consistent improvement over the baseline (QWK: 27.87%, F1: 38.02%), achieving more predictable and interpretable results. These results suggest the controlled approach to generating learner errors better captures the actual distribution of errors across proficiency levels. In future work, we plan to expand the dataset by incorporating real essay samples, which will enhance the linguistic and error profiles for each CEFR level, thereby improving the proposed data augmentation model. In addition, we will deploy the AES system as part of an interactive educational writing assistant tool to help students enhance their Arabic writing with immediate feedback on errors and CEFR scores.
8 Limitations
Despite the effectiveness of our hybrid Arabic AES framework, we note several limitations related to the quality of generated Arabic essays, error injection accuracy, and the generalization of the AES model. The lack of A1 and C2 essays in ZAEBUC means that there is no gold reference data for these levels, which may impact both linguistic and error profiles, affecting the accuracy of GPT-generated essays. Furthermore, different biases exist in both ZAEBUC and the GPT outputs: ZAEBUC overrepresents B-level essays, skewing model training; and GPT-4o tends to favor Arabic culture themes, thereby limiting diversity in the generated texts, and usability in other contexts. In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles. Another limitation is the model’s ability to generalize across various domains and question types. The AES system may struggle with broader writing tasks and alternative prompts since the dataset and augmentation methods focus on predefined prompts. Relying solely on CEFR as a holistic scoring method limits interpretability. Enhancing the dataset with multi-trait annotations, such as coherence, argumentation, and organization, could improve scoring accuracy and feedback quality. Moreover, better-controlled GPT prompting could refine the quality and diversity of generated essays, reducing biases and improving alignment with real learner writing patterns.
9 Ethical Considerations
While Arabic AES systems provide significant support in assessing Arabic learners’ writing proficiency, it is essential to highlight the ethical implications of their use. Automatic assessment and scoring may lead to misjudgments that could distress learners and students, especially if their work is incorrectly evaluated. AES tools should serve as an educational assistive technology, complementing the teacher’s judgment, not replacing it in educational settings.
Appendix A GPT-4o Error injection algorithm
Figure 6: GPT-4o Error injection and verification algorithm